# On the non-Markovian quantum control dynamics

### v2

@@ III-A $\delta=0$, $g_{n}\neq 0$
where the non-Markovian interaction between the atom and the environment can influence the integration kernel in Eq. (38), and further influence the convergence rate and steady value of $\tilde{X}$. The dynamics of the equation with the format of Eq. (38) has been introduced in Ref. [71]. Generalized from Theorem 2 in Ref. [71] (Chapter 2, Page 36), we have the following proposition.
Proposition 8.
Provided that $\|\textbf{P}(t)\|\leq c_{1}$ for $t\geq t_{0}$, where $c_{1}$ is determined by $\max(L_{n}^{(1)})$, it follows that $\lim_{t\rightarrow\infty}\tilde{X}(t)=0$ in Eq. (38).
Proof.
The proof is similar to that in Ref. [71] and is thus omitted due to page limitations. ∎
Proposition 8 means that when the interaction between the atom and environment converges to a Markovian regime, $\tilde{X}(t)$ will converge to zero. Besides, Proposition 8 is further generalized to linear time-varying system in Ref. [71], and this generalization will be used in the following analysis.
III-B $\delta\neq 0$, $g_{n}\neq 0$
In the following, we consider the most general case by separating $X(t)$ by its real and imaginary parts. Considering that in Eq. (9), Eq. (9a) and Eq. (9d) are always real, Eq. (9b) and Eq. (9c) are conjugate complex values, then we rewrite the state vector in a simplified format as
$\mathbb{X}(t)=\left[\mathbb{X}_{1}(t),\mathbb{X}_{2}(t),\cdots,\mathbb{X}_{N-1}(t),a^{R}\right]^{T},$
@@ III-B $\delta\neq 0$, $g_{n}\neq 0$
which is a real-value equation. We take the $n$-th energy level as an example, $\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{X}_{n}(t)=\mathbb{A}_{n}(t)\mathbb{X}_{n}(t)+\mathbb{D}_{n}(t)a^{R}$, and
$\mathbb{A}_{n}(t)=\begin{bmatrix}-2R_{f}(t)&2g_{n}\sin\delta t&2g_{n}\cos\delta t\\ -g_{n}\sin\delta t&-R_{f}(t)&I_{f}(t)-\delta\\ -g_{n}\cos\delta t&\delta-I_{f}(t)&-R_{f}(t)\\ \end{bmatrix},$ (40)
$\mathbb{C}_{n}(t)$ $=\begin{bmatrix}0&-2g_{n}\sin\delta t&-2g_{n}\cos\delta t\end{bmatrix},$ (41)
$\mathbb{D}_{n}(t)$ $=\begin{bmatrix}0\\ -g_{n}\sin\delta t\\ -g_{n}\cos\delta t\end{bmatrix},$ (42)
where we take $F(t)|\kappa_{n}|^{2}\triangleq R_{f}(t)+iI_{f}(t)$, $F^{*}(t)|\kappa_{n}|^{2}\triangleq R_{f}(t)-iI_{f}(t)$ for simplification.
In summary, when there are detunings between the atom and cavity, the dynamics of atomic states and the photon number in the cavity will always be linear and time-varying.
Upon this, when the interaction between the atom and the environment becomes asymptotically Markovian, we denote $\lim_{t\rightarrow\infty}R_{f}(t)=\bar{R}_{f}$ and $\lim_{t\rightarrow\infty}I_{f}(t)=\bar{I}_{f}$. Then Eq. (40) becomes
$\mathbb{A}_{n}(t)=$ $\begin{bmatrix}-2\bar{R}_{f}&2g_{n}\sin\delta t&2g_{n}\cos\delta t\\ -g_{n}\sin\delta t&-\bar{R}_{f}&\bar{I}_{f}-\delta\\ -g_{n}\cos\delta t&\delta-\bar{I}_{f}&-\bar{R}_{f}\\ \end{bmatrix}$ (43)
$+\begin{bmatrix}2\bar{R}_{f}(t)-2R_{f}(t)&0&0\\ 0&\bar{R}_{f}-R_{f}(t)&I_{f}(t)-\bar{I}_{f}\\ 0&\bar{I}_{f}-I_{f}(t)&\bar{R}_{f}-R_{f}(t)\\ \end{bmatrix}$
@@
where $\bar{\mathbb{A}}_{n}$ represents the first matrix in Eq. (43) and $\tilde{\mathbb{A}}_{n}$ represents the second matrix. Obviously, $\bar{\mathbb{A}}_{n}(t)$ is periodic with $\bar{\mathbb{A}}_{n}(t)=\bar{\mathbb{A}}_{n}(t+T_{\delta}),$ where $T_{\delta}=\frac{2\pi}{\delta}$, and the norm of $\tilde{\mathbb{A}}_{n}(t)$ finally converges to zero.
To clarify the stability of the above linear time-varying system, we first introduce the following definition and propositions.
Definition 4 ([72]).
$\|\Phi\left(t,\tau\right)\|\leq Ke^{-\tilde{\alpha}\left(t-\tau\right)},for~{}t_{0}\leq\tau<t<\infty.$
Definition 5 ([72]).
$\mu[A(t)]=\lim_{h\rightarrow 0^{+}}\frac{\|I_{n}+hA(t)\|-1}{h},$
and
$\Pi^{+}(t)\triangleq\int_{t_{0}}^{t}\mu[A(\tau)]\mathrm{d}\tau.$ (44)
As a combination of the conclusion in Ref. [72] on the relationship between $\Pi^{+}(t)$ and the stability of a periodic linear time-varying system, and the generalization of Proposition 8 to the linear time-varying system in Ref. [71], we can derive the following stability proposition for the linear time-varying system in Eq. (39) with the time-varying matrix in Eq. (43).
Proposition 9.
For the linear time-varying system with the matrix represented as a combination of time-varying periodic and time-varying converging matrices, as given in Eq. (43), the solutions approach zero when
@@
The proof is based on two separated parts as discussed in Refs. [72, 71]. According to Ref. [72], when (a) is satisfied, the linear time-varying system $\dot{\hat{\mathbb{X}}}=\bar{\mathbb{A}}_{n}(t)\hat{\mathbb{X}}$ is uniformly exponential stable, and $\lim_{t\rightarrow\infty}\hat{\mathbb{X}}=0$. Based on this, when (b) is satisfied, the state vector in Eq. (39) approaches zero, as shown in Ref. [71] (Chapter 2, Theorem 4). ∎
III-C With external drives
@@
Then the evolution of the mean values of operators reads
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{+}_{n}a\rangle=$ $-i\delta\langle\sigma^{+}_{n}a\rangle-ig_{n}e^{-i\delta t}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle+ig_{n}e^{-i\delta t}\langle a^{{\dagger}}a\rangle$ (46)
$-F^{*}(t)|\kappa_{n}|^{2}\langle\sigma^{+}_{n}a\rangle+\mathcal{E}\left\langle\sigma^{+}_{n}\right\rangle,$
$\frac{\mathrm{d}}{\mathrm{d}t}\langle a^{{\dagger}}a\rangle=$ $i\sum_{n}g_{n}\left(e^{i\delta t}\langle\sigma^{+}_{n}a\rangle-e^{-i\delta t}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle\right)$ (47)
$+\mathcal{E}\langle a+a^{{\dagger}}\rangle,$
$\frac{\mathrm{d}}{\mathrm{d}t}\left\langle\sigma^{-}_{n}\right\rangle=-ig_{n}e^{i\delta t}\langle a\rangle-F(t)|\kappa_{n}|^{2}\left\langle\sigma^{-}_{n}\right\rangle.$ (48)
Obviously, the drive field directly influences the evolution of the operators containing the cavity operator. We further denote $R_{a}=\left\langle a+a^{{\dagger}}\right\rangle$, and $I_{a}=i\left\langle a-a^{{\dagger}}\right\rangle$ with real values, then
$\frac{\mathrm{d}}{\mathrm{d}t}R_{a}$ $=-i\sum_{n}g_{n}e^{-i\delta t}\left\langle\sigma^{-}_{n}\right\rangle+i\sum_{n}g_{n}e^{i\delta t}\left\langle\sigma^{+}_{n}\right\rangle+2\mathcal{E}$ (49)
$=2\sum_{n=1}^{N-1}g_{n}\left(s_{n}^{I}\cos\delta t-s_{n}^{R}\sin\delta t\right)+2\mathcal{E},$
$\frac{\mathrm{d}}{\mathrm{d}t}I_{a}$ $=\sum_{n}g_{n}e^{-i\delta t}\left\langle\sigma^{-}_{n}\right\rangle+\sum_{n}g_{n}e^{i\delta t}\left\langle\sigma^{+}_{n}\right\rangle$ (50)
$=2\sum_{n=1}^{N-1}g_{n}\left(s_{n}^{R}\cos\delta t+s_{n}^{I}\sin\delta t\right).$
where $s_{n}^{R}$ and $s_{n}^{I}$ represent the real and imaginary parts of $\left\langle\sigma^{-}_{n}\right\rangle$, respectively.
Generalized from the state vector $\mathbb{X}(t)$ in Eq. (39), we define another real-valued vector $\tilde{\mathbb{X}}$ as
$\tilde{\mathbb{X}}(t)=\left[\tilde{\mathbb{X}}_{1}(t),\tilde{\mathbb{X}}_{2}(t),\cdots,\tilde{\mathbb{X}}_{N-1}(t),a^{R},R_{a},I_{a}\right]^{T},$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{\mathbb{X}}$ $=\begin{bmatrix}\mathbb{B}_{1}(t)&0&\cdots&0&\tilde{\mathbb{D}}_{1}(t)\\ 0&\mathbb{B}_{2}(t)&\cdots&0&\tilde{\mathbb{D}}_{2}(t)\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\mathbb{B}_{N-1}(t)&\tilde{\mathbb{D}}_{N-1}(t)\\ \tilde{\mathbb{C}}_{1}(t)&\tilde{\mathbb{C}}_{2}(t)&\cdots&\tilde{\mathbb{C}}_{N-1}(t)&\tilde{\mathbb{P}}\end{bmatrix}\tilde{\mathbb{X}}+2\mathcal{E}\begin{bmatrix}0\\ 0\\ \vdots\\ 1\\ 0\end{bmatrix}$ (51)
$\triangleq\mathbb{B}(t)\tilde{\mathbb{X}}+2\mathcal{E}\begin{bmatrix}0\\ 0\\ \vdots\\ 1\\ 0\end{bmatrix},$
@@
$\mathbb{B}_{n}(t)=\begin{bmatrix}-2R_{f}(t)&2g_{n}\sin\delta t&2g_{n}\cos\delta t&0&0\\ -g_{n}\sin\delta t&-R_{f}(t)&I_{f}(t)-\delta&\mathcal{E}&0\\ -g_{n}\cos\delta t&\delta-I_{f}(t)&-R_{f}(t)&0&-\mathcal{E}\\ 0&0&0&-R_{f}(t)|\kappa_{n}|^{2}&I_{f}(t)|\kappa_{n}|^{2}\\ 0&0&0&-I_{f}(t)|\kappa_{n}|^{2}&-R_{f}(t)|\kappa_{n}|^{2}\end{bmatrix},$ (52)
~$\tilde{\mathbb{C}}_{n}(t)$ $=\begin{bmatrix}0&-2g_{n}\sin\delta t&-2g_{n}\cos\delta t&0&0\\ 0&0&0&-2g_{n}\sin\delta t&2g_{n}\cos\delta t\\ 0&0&0&2g_{n}\cos\delta t&2g_{n}\sin\delta t\end{bmatrix},$ (53)
$\tilde{\mathbb{D}}_{n}(t)$ $=\begin{bmatrix}0&0&0\\ -g_{n}\sin\delta t&0&0\\ -g_{n}\cos\delta t&0&0\\ 0&\frac{g_{n}\sin\delta t}{2}&-\frac{g_{n}\cos\delta t}{2}\\ 0&-\frac{g_{n}\cos\delta t}{2}&\frac{g_{n}\sin\delta t}{2}\end{bmatrix},$ (54)
@@
We take the two-level atom as an example, to clarify how the dynamics can be influenced by the non-Markovian environment, coupling strength between the atom and cavity, and the external drive applied on the cavity, which are compared in Fig. 4. In all the four simulations, $\omega_{a}=50$, $\delta=10$, and $\Omega=45$, $\kappa_{1}=4$, $\chi=1$, $\gamma=1$ for the non-Markvoian parameter settings, which is the same as that in solid lines of Fig. 3. In Fig. 4(a-b), $\mathcal{E}=0.02$, and in (c-d), $\mathcal{E}=0.2$. When the drive applied on the cavity is weak, as in (a-b), the atom can decay to the ground state and the cavity is finally almost empty, because of the non-Markovian Lindblad components. However, when $\mathcal{E}$ is larger and the coupling between the atom and cavity is not so strong, then there can be multiple photons in the cavity, as in (c). When both $\mathcal{E}$ and $g_{1}$ are larger, the exchanging between the atom and cavity is faster, and the atom and cavity states are both oscillating apart from decaying to the environment, thus the population of the excited atom and photon numbers in the cavity are not exponentially stable, as shown in (d).
Figure 4: The non-Markovian dynamics of quantum system with one two-level atom in the cavity with different drives.
@@ IV Quantum measurement feedback control in the non-Markovian cavity-QED system
$I_{c}(t)$ $=\sqrt{2}\langle x\rangle+\frac{1}{\sqrt{\eta}}\xi(t),$ (56)
where $x=\left(a+a^{{\dagger}}\right)/\sqrt{2}$, $\xi(t)$ is the measurement-induced time-independent random term satisfies that $\langle\xi(t)\rangle=0$ and $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$, and $\eta$ is the measurement detection efficiency. Then the feedback Hamiltonian reads
@@
Assume that the detection efficiency of the measurement apparatus is ideal as $\eta=1$.
The Markovian stochastic master equation with measurement feedback can be rewritten as [34, 74]
$\mathrm{d}\rho$ $=-i[H,\rho(t)]\mathrm{d}t+\kappa\mathcal{L}_{a}\left[\rho(t)\right]\mathrm{d}t-ig_{f}\left[G,a\rho(t)+\rho(t)a^{{\dagger}}\right]\mathrm{d}t$ (58)
$~{}~{}~{}~{}-\frac{1}{2}g_{f}^{2}\left[G,\left[G,\rho(t)\right]\right]\mathrm{d}t+\mathcal{H}[a]\rho(t)dW-ig_{f}\left[G,\rho(t)\right]\mathrm{d}W,$
where $\mathcal{L}_{a}$ represents the decaying of the cavity with the rate $\kappa$, $\mathcal{L}_{O}\left[\rho\right]=O\rho O^{{\dagger}}-\frac{1}{2}\rho O^{{\dagger}}O-\frac{1}{2}O^{{\dagger}}O\rho$ for an arbitrary operator $O$, $\mathcal{H}[O]\rho=O\rho(t)+\rho(t)O^{{\dagger}}-\rm{Tr}\left[\right(O+O^{{\dagger}}\left)\rho(t)\right]\rho(t)$, $g_{f}$ is the feedback strength, $\mathrm{d}W(t)=\xi(t)\mathrm{d}t$ is a Wiener process representing the Homodyne detection noise in the quantum measurement, $E\left[\mathrm{d}W(t)^{2}\right]=\mathrm{d}t$, and $\xi(t)$ can be approximated with the white noise [75].
@@
Generalized from Eq. (58), the stochastic quantum measurement feedback control equation when there are non-Markovian interactions between the atom and environment reads
$\mathrm{d}\rho=$$\left\{-i[\bar{H}+H_{D},\rho(t)]+\left(F^{*}(t)+F(t)\right)L\rho(t)L^{{\dagger}}-F(t)L^{{\dagger}}L\rho(t)\right.$ (59)
$\left.-F^{*}(t)\rho(t)L^{{\dagger}}L+\kappa\mathcal{L}_{a}\left[\rho(t)\right]\right\}\mathrm{d}t-ig_{f}\left[G,a\rho(t)+\rho(t)a^{{\dagger}}\right]\mathrm{d}t$
$-\frac{1}{2}g_{f}^{2}\left[G,\left[G,\rho(t)\right]\right]\mathrm{d}t+\mathcal{H}[a]\rho(t)dW-ig_{f}\left[G,\rho(t)\right]\mathrm{d}W,$
where $H_{D}$ represents external drives applied upon the cavity, $F(t)$ is for the non-Markovian interaction between the atom and the environment, and the other components have the same meaning as in Eq. (58).
In the following, we study two circumstances. One is that the feedback operator $G$ is an atomic operator, the other is that $G$ is a cavity operator.
IV-A Feedback control with atomic operators
The feedback operator for the $N$-level system can be represented as
IV-A1 Example
When $N=2$, we choose $G=\left(|0\rangle\langle 1|+|1\rangle\langle 0|\right)$, and assume initially the atom is excited. The atomic state can be evaluated with the vector of the operators as $S=\left[\langle a\rangle,\langle\sigma_{-}\rangle,\langle\sigma_{z}\rangle\right]^{T}\triangleq\left[\bar{\alpha},\bar{s},\bar{w}\right]^{T}$, where $\sigma_{-}=|0\rangle\langle 1|$ and $\sigma_{z}=|1\rangle\langle 1|-|0\rangle\langle 0|$. When the external drive is $H_{D}=-i\mathcal{E}\left(a-a^{{\dagger}}\right)$, the dynamics of the cavity or atomic operator reads
$\frac{\mathrm{d}}{\mathrm{d}t}\bar{\alpha}=-ig_{1}e^{-i\delta t}\bar{s}-\frac{\kappa}{2}\bar{\alpha}+\mathcal{E},$ (61a)
$\frac{\mathrm{d}}{\mathrm{d}t}\bar{s}=ig_{1}e^{i\delta t}\bar{\alpha}\bar{w}-F(t)|\kappa_{1}|^{2}\bar{s}+ig_{f}\left(\bar{\alpha}+\bar{\alpha}^{*}\right)\bar{w}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}-\left(F(t)+F^{*}(t)\right)|\kappa_{1}|^{2}\left(\bar{w}+1\right)-2ig_{f}\left(\bar{s}^{*}-\bar{s}\right)\left(\bar{\alpha}+\bar{\alpha}^{*}\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}-2g_{f}^{2}\bar{w}-2ig_{f}\left(\bar{s}^{*}-\bar{s}\right)\xi(t),$ (61c)
with the initially normalization condition that $\bar{w}^{2}+4\left|\bar{s}\right|^{2}=1$ [76], and the stochastic component with higher order amplitude is omitted in Eq. (61a). We consider the simplified steady states satisfying that $\dot{\bar{\alpha}}=\dot{\bar{s}}=\dot{\bar{w}}=0$ after averaging the measurement noise $\xi(t)=0$, and take $\delta=0$ for simplification.
Figure 5: Measurement feedback control for one two-level atom in the cavity.
The performance for different non-Markovian parameter settings is compared in Fig. 5, where we take $\kappa_{1}=1$, $\chi=0.5$, $\omega_{a}=50$, $\Omega=45$, $\mathcal{E}=0.1$, $g_{1}=0.2$, $\delta=0$, $\kappa=0.2$ and $\mathrm{d}t=0.01$. It can be seen that the measurement feedback control can enhance the probability that the atom is excited, both when the interface between atom and environment is Markovian and non-Markovian.
IV-B Feedback control with cavity operators
Additionally, the feedback operator can be designed according to the cavity operator as [77]

### v3

@@ III Time-varying linear quantum control dynamics
where the non-Markovian interaction between the atom and the environment can influence the integration kernel in Eq. (38), and further influence the convergence rate and steady value of $\tilde{X}$. The dynamics of the equation with the format of Eq. (38) has been introduced in Ref. [58]. Generalized from Theorem 2 in Ref. [58] (Chapter 2, Page 36), we have the following corollary.
Corollary 1.
Provided that $\|\textbf{P}(t)\|\leq c_{1}$ for $t\geq t_{0}$ in Eq. (38), where $c_{1}$ is determined by $\max(L_{n}^{(1)})$, it follows that $\lim_{t\rightarrow\infty}\tilde{X}(t)=0$ in Eq. (38).
Corollary 1 means that when the interaction between the atom and environment converges to a Markovian regime, $\tilde{X}(t)$ will converge to zero. Besides, Corollary 1 is further generalized to linear time-varying system in Ref. [58], and this generalization will be used in the following analysis.
III-A $\Delta_{n}\neq 0$, $g_{n}\neq 0$
In the following, we consider the most general case by separating $X(t)$ into its real and imaginary parts. Considering that in Eq. (9), the LHS of Eq. (9a) and Eq. (9d) are always real, Eq. (9b) and Eq. (9c) are conjugate complex values, then we rewrite the state vector in a simplified format as
$\mathbb{X}(t)=\left[\mathbb{X}_{1}(t),\mathbb{X}_{2}(t),\cdots,\mathbb{X}_{N-1}(t),a^{R}\right]^{\rm T},$
@@ III-A $\Delta_{n}\neq 0$, $g_{n}\neq 0$
which is a real-value equation. We take the $n$-th energy level as an example, denote $\hat{g}_{n}=g_{n}\sin\left(\Delta_{n}t\right)$, $\breve{g}_{n}=g_{n}\cos\left(\Delta_{n}t\right)$, and $\dot{\mathbb{X}}_{n}(t)=\mathbb{A}_{n}(t)\mathbb{X}_{n}(t)+\mathbb{D}_{n}(t)a^{R}$, then
$\mathbb{A}_{n}(t)=\begin{bmatrix}-2\mathcal{R}_{n}(t)&2\hat{g}_{n}&2\breve{g}_{n}\\ -\hat{g}_{n}&-\mathcal{R}_{n}(t)&\mathcal{I}_{n}(t)-\Delta_{n}\\ -\breve{g}_{n}&\Delta_{n}-\mathcal{I}_{n}(t)&-\mathcal{R}_{n}(t)\\ \end{bmatrix},$ (40)
$\mathbb{C}_{n}(t)$ $=\begin{bmatrix}0&-2\hat{g}_{n}&-2\breve{g}_{n}\end{bmatrix},$ (41)
$\mathbb{D}_{n}(t)$ $=\begin{bmatrix}0&-\hat{g}_{n}&-\breve{g}_{n}\end{bmatrix}^{\rm T},$ (42)
where we take $F_{n}(t)\kappa_{n}\triangleq\mathcal{R}_{n}(t)+i\mathcal{I}_{n}(t)$, $F_{n}^{*}(t)\kappa_{n}\triangleq\mathcal{R}_{n}(t)-i\mathcal{I}_{n}(t)$ for simplification.
Upon this, when the interaction between the atom and the environment becomes asymptotically Markovian, we denote $\lim_{t\rightarrow\infty}\mathcal{R}_{n}(t)=\bar{\mathcal{R}}_{n}$ and $\lim_{t\rightarrow\infty}\mathcal{I}_{n}(t)=\bar{\mathcal{I}}_{n}$. Then Eq. (40) becomes
$\mathbb{A}_{n}(t)=\begin{bmatrix}-2\bar{\mathcal{R}}_{n}&2\hat{g}_{n}&2\breve{g}_{n}\\ -\hat{g}_{n}&-\bar{\mathcal{R}}_{n}&\bar{\mathcal{I}}_{n}-\Delta_{n}\\ -\breve{g}_{n}&\Delta_{n}-\bar{\mathcal{I}}_{n}&-\bar{\mathcal{R}}_{n}\\ \end{bmatrix}$ (43)
$+\begin{bmatrix}2\bar{\mathcal{R}}_{n}-2\mathcal{R}_{n}(t)&0&0\\ 0&\bar{\mathcal{R}}_{n}-\mathcal{R}_{n}(t)&\mathcal{I}_{n}(t)-\bar{\mathcal{I}}_{n}\\ 0&\bar{\mathcal{I}}_{n}-\mathcal{I}_{n}(t)&\bar{\mathcal{R}}_{n}-\mathcal{R}_{n}(t)\\ \end{bmatrix}$
@@
where $\bar{\mathbb{A}}_{n}$ represents the first matrix in Eq. (43) and $\tilde{\mathbb{A}}_{n}$ represents the second matrix. Obviously, $\bar{\mathbb{A}}_{n}(t)$ is periodic with $\bar{\mathbb{A}}_{n}(t)=\bar{\mathbb{A}}_{n}\left(t+2\pi/\Delta_{n}\right)$, and the norm of $\tilde{\mathbb{A}}_{n}(t)$ finally converges to zero.
As a combination of the conclusion in Ref. [59] on the relationship between $\Pi^{+}(t)$ (see definition in [59] or Appendix C), the stability of a periodic linear time-varying system, and the generalization of Corollary 1 to the linear time-varying system in Ref. [58], we can derive the following corollary for the linear time-varying system in Eq. (39) with the time-varying matrix in Eq. (43).
Corollary 2.
For the linear time-varying system (39) with the matrix represented as a combination of time-varying periodic and time-varying converging matrices, as given in Eq. (43), the solutions approach zero when
@@
See Appendix D. ∎
III-B With external drives
@@
We denote $R_{a}=\left\langle a+a^{{\dagger}}\right\rangle$, and $I_{a}=i\left\langle a-a^{{\dagger}}\right\rangle$ with real values, and $s_{n}^{R}$ and $s_{n}^{I}$ represent the real and imaginary parts of $\left\langle\sigma^{-}_{n}\right\rangle$, respectively. Then generalized from the state vector $\mathbb{X}(t)$ in Eq. (39), we define another real-valued vector $\tilde{\mathbb{X}}$ as
$\tilde{\mathbb{X}}(t)=\left[\tilde{\mathbb{X}}_{1}(t),\tilde{\mathbb{X}}_{2}(t),\cdots,\tilde{\mathbb{X}}_{N-1}(t),a^{R},R_{a},I_{a}\right]^{\rm T},$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{\mathbb{X}}$ $=\mathbb{B}(t)\tilde{\mathbb{X}}+2\mathcal{E}\begin{bmatrix}0,0,\cdots,1,0\end{bmatrix}^{\rm T},$ (45)
@@
$\mathbb{B}(t)=\begin{bmatrix}\mathbb{B}_{1}(t)&0&\cdots&0&\tilde{\mathbb{D}}_{1}(t)\\ 0&\mathbb{B}_{2}(t)&\cdots&0&\tilde{\mathbb{D}}_{2}(t)\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\mathbb{B}_{N-1}(t)&\tilde{\mathbb{D}}_{N-1}(t)\\ \tilde{\mathbb{C}}_{1}(t)&\tilde{\mathbb{C}}_{2}(t)&\cdots&\tilde{\mathbb{C}}_{N-1}(t)&\tilde{\mathbb{P}}\end{bmatrix},$ (46)
$\mathbb{B}_{n}=$ (47)
$\begin{bmatrix}-2\mathcal{R}_{n}&2\hat{g}_{n}&2\breve{g}_{n}&0&0\\ -\hat{g}_{n}&-\mathcal{R}_{n}&\mathcal{I}_{n}-\Delta_{n}&\mathcal{E}&0\\ -\breve{g}_{n}&\Delta_{n}-\mathcal{I}_{n}&-\mathcal{R}_{n}&0&-\mathcal{E}\\ 0&0&0&-\mathcal{R}_{n}&\mathcal{I}_{n}\\ 0&0&0&-\mathcal{I}_{n}&-\mathcal{R}_{n}\end{bmatrix},$
~$\tilde{\mathbb{C}}_{n}$ $=\begin{bmatrix}0&-2\hat{g}_{n}&-2\breve{g}_{n}&0&0\\ 0&0&0&-2\hat{g}_{n}&2\breve{g}_{n}\\ 0&0&0&2\breve{g}_{n}&2\hat{g}_{n}\end{bmatrix},$ (48)
$\tilde{\mathbb{D}}_{n}$ $=\begin{bmatrix}0&0&0\\ -\hat{g}_{n}&0&0\\ -\breve{g}_{n}&0&0\\ 0&\frac{\hat{g}_{n}}{2}&-\frac{\breve{g}_{n}}{2}\\ 0&-\frac{\breve{g}_{n}}{2}&\frac{\hat{g}_{n}}{2}\end{bmatrix},$ (49)
@@
Figure 4: The non-Markovian dynamics of quantum system with one two-level atom in the cavity with applied drives.
We take the two-level atom as an example to illustrate how the dynamics can be influenced by the non-Markovian environment, coupling strength between the atom and cavity, and the external drive applied on the cavity, which are compared in Fig. 4. In all the simulations, $\omega_{1}=37.7$GHz, $\kappa_{1}=0.31$GHz, the time step for simulations and the non-Markovian environmental parameter settings are the same as in Fig. 3. In Fig. 4(a), the cavity is resonant with the atom, and we compared how the dynamics is influenced by the coupling strength $g_{1}$ when $\mathcal{E}=10$MHz. It can be seen that when the cavity is driven, there can be multi-photon states in the cavity when the coupling strength between the atom and cavity is weak. In Fig. 4(b), the atom is initially in a superposition state, $g_{1}=20$MHz, $\Delta_{1}=-20$MHz, there can also be multi-photon states when the cavity is driven.
@@ IV Quantum measurement feedback control in the non-Markovian cavity-QED system
For general cavity-QED systems, the decaying process of both the atom and cavity to the environment can be regarded as Markovian or non-Markovian. The non-Markovian interaction between the cavity and environment can be generalized from Appendixes A and B by replacing the atomic operator with the cavity operator. We further clarify the quantum measurement feedback control in this section based on the above generalization.
@@
$I_{c}(t)$ $=\sqrt{2}\langle x\rangle+\xi(t)/\sqrt{\eta},$ (51)
where $x=\left(a+a^{{\dagger}}\right)/\sqrt{2}$, $\xi(t)$ is the measurement-induced random term satisfies that $\langle\xi(t)\rangle=0$ and $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$, and $\eta$ is the measurement detection efficiency. The feedback Hamiltonian reads
@@
where $G=G^{{\dagger}}$ is the feedback operator. We assume that the detection efficiency of the measurement apparatus is ideal as $\eta=1$.
@@
Generalized from Eq. (6) and the Markovian stochastic master equation (SME) with measurement feedback control in Refs. [19, 61], the measurement feedback based SME when there are non-Markovian interactions between the quantum system and environment reads [60]
$\mathrm{d}\rho=-i\left[\bar{H}+H_{D},\rho\right]\mathrm{d}t+\sum_{n=1}^{N-1}\left[\left(F_{n}+F_{n}^{*}\right)L_{n}\rho L_{n}^{{\dagger}}\right.$ (53)
$\left.-\left(F_{n}L_{n}^{{\dagger}}L_{n}\rho+F_{n}^{*}\rho L_{n}^{{\dagger}}L_{n}\right)\right]\mathrm{d}t+\kappa\left[F_{a}+F_{a}^{*}\right]a\rho a^{{\dagger}}\mathrm{d}t$
$-\kappa\left[F_{a}a^{{\dagger}}a\rho+F_{a}^{*}\rho a^{{\dagger}}a\right]\mathrm{d}t-ig_{f}\left[G,a\rho+\rho a^{{\dagger}}\right]\mathrm{d}t$
$-\frac{g_{f}^{2}}{2}\left[G,\left[G,\rho\right]\right]\mathrm{d}t+\mathcal{H}[a]\rho\mathrm{d}W-ig_{f}\left[G,\rho\right]\mathrm{d}W,$
where $H_{D}$ represents external drives applied upon the cavity, $F_{a}$ is for the non-Markovian interaction between the cavity and the environment, and is governed by the similar dynamics as $F_{n}$ in Eq. (80) in Appendix A by respectively replacing the parameters $\omega_{n}$, $\kappa_{n}$ and $\chi_{n}$ with $\omega_{c}$, $\kappa_{c}$ for the coupling strength between the cavity and environment, and $\chi_{a}$ in analog to $\chi_{n}$ for the environment parameter setting, $\kappa$ represents a constant decay rate of the cavity to the environment, $\mathcal{H}[O]\rho=O\rho+\rho O^{{\dagger}}-\rm{Tr}\left[\right(O+O^{{\dagger}}\left)\rho\right]\rho$ for an arbitrary operator $O$, $g_{f}$ is the feedback strength, $\mathrm{d}W(t)=\xi(t)\mathrm{d}t$ is for a Wiener process representing the Homodyne detection noise in the quantum measurement, $E\left[\mathrm{d}W(t)^{2}\right]=\mathrm{d}t$, $\xi(t)$ can be approximated with the white noise [62], and the meaning of the other components is the same as Eq. (6).
In the following, we firstly clarify the circumstance that the feedback operator $G$ is a cavity operator, then illustrate the circumstance that $G$ is an atomic operator with numerical simulations.
IV-A Feedback control with cavity operators
In the following, we assume that the interaction between the cavity and environment is Markovian, by taking $F_{a}\equiv\kappa/2$. Additionally, the feedback operator can be designed as [63]